OpenStudy (anonymous):

Solve -15(y – 2) – 3 = -5y –11(3y + 7). Please show all of your work

OpenStudy (anonymous):

23y=-104 y=-104/23

OpenStudy (anonymous):

-3-15 (-2+y) = -5 y-11 (7+3 y) -3-15 (-2)-15 y = -5 y-11 7-11 3 y -3-2 (-1) 15-15 y = -5 y-7 11-3 11 y -3-30 (-1)-15 y = -5 y-7 11-3 11 y -3+30-15 y = -5 y-77-33 y = 27-15 y = -38 y-77 -23y = 104 y = -104/23

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just making sure i did it right

OpenStudy (anonymous):

thanks